# Learning Collective Variables for Enhanced Sampling from BioEmu with Time-Lagged Generation

### v2

@@ 5 Conclusion
We present BioEmu-CV, a lightweight framework that learns collective variables (CVs) from time-lagged conditioning signals on a frozen BioEmu foundation model. Given the current structure, we derive a low-dimensional CV and inject it into BioEmu’s single representation; where only this conditioning path is trained with a denoising score-matching objective, leaving the backbone untouched. We systematically evaluate MLCVs on three fast-folding proteins with two downstream tasks, estimating free energy differences with OPES simulations and transition path sampling with CV-steered MD. Our experiments that BioEmu-CV successfully identifies important macroscopic movements related to the slow degree of freedom, along with qualitative evaluation on state discrimination.
@@
All protein trajectory data used in this study is publicly available from the D. E. Shaw Research (DESRES) database. The detailed description of the dataset can be found in lindorff2011fast. We provided brief explanation of the dataset in Section A.2. Implementation details for our method and all baselines, including model and training hyperparameters, are specified in Section B.2. Furthermore, the experimental setups for our downstream evaluation tasks are provided in the appendix; Section B.3 contains the configurations for the on-the-fly probability enhanced sampling (OPES) simulations, and Section B.4 outlines the setup for the steered molecular dynamics (SMD) simulations. For complete reproduction, we provide our code.
The Use of Large Language Models (LLMs)
LLMS were only used to aid and polish writing, and has not played a significant role in research ideation and/or writing.
@@ A.3 Time-lagged independent component analysis
Appendix B Simulation and evaluation details
B.1 Qualitative verification
@@ B.2 Baseline details
B.3 OPES simulation details
OPES configurations. We use configurations as in Table 9 for PLUMED, differing only in temperatures. A PACE of 500 steps is applied, i.e., 1 picoseconds. We use the same SIGMA value 0.05 and BARRIER value 30 for all systems. Temperatures are identical to the reference simulation. The first 100 ns are discarded as equilibration time with 50 ns window unit time steps for Figures 2 and 10.
Protein PACE SIGMA BARRIER (kJ/mol) Temperature (K)
Chignolin 500 0.05 30 340
Trp-cage (TC10b) 500 0.05 30 290
BBA 500 0.05 30 325
B.4 Steered MD simulation details
Appendix C Additional results
C.1 OPES simulation results
C.2 State discrimination
C.3 Steered MD visualization
C.4 VAMP score
Method Chignolin Trp-cage BBA
VAMP-1 VAMP-2 VAMP-E VAMP-1 VAMP-2 VAMP-E VAMP-1 VAMP-2 VAMP-E
TAE 1.9686 1.9381 1.9381 1.8787 1.7722 1.7722 1.8902 1.7925 1.7925
VDE 1.9850 1.9702 1.9702 1.9947 1.9894 1.9894 1.8469 1.7172 1.7172
C.5 Simple baselines
Experiment setup. Here, we show that simple baselines, e.g., PCA and TICA, on $C_{\alpha}$-wise distance fail for Trp-cage. We use the first principal component and the first time-lagged principal component as CVs. All other details are identical; the training data is for the baseline and BioEmu-CV, normalization to the range [-1, 1] over the whole DESRES trajectory data.
@@ C.5 Simple baselines
C.6 Capturing multi-pathways in Alanine Dipeptide
Here, we test CVs from time-lagged generation for capturing multi-pathways in Alanine Dipeptide.
Alanine Dipeptide. Alanine Dipeptide is a well-studied molecule consisting of 22 atoms with two optimal CVs: the backbone dihedral angles $\phi$ and $\psi$. We use two meta-stable states $C5$ and $C7_{ax}$, each located at (2.49, 2.67)(-2.49,\ 2.67) and (1.02,0.70)(1.02,\ -0.70) in $(\phi,\psi)$. Typically, two pathways are known to pass near each saddle point in the energy barrier at $\phi=0$ (holdijk2023stochastic; seong2024transition). Training data were randomly extracted from 10 ns molecular dynamics trajectories with amber99sbildn force field and tip3p implicit solvent by OpenMM (eastman2023openmm) for a fair comparison, with five trajectories each initialized in the $C5$ and $C7_{ax}$ meta-stable state. We train from scratch with time-lagged conditioning, since the released weights were obtained from a different force field. Among data pairs $(x_{t},x_{t+\tau})$, no transition events exists.
Experiment setup. We use the Transferable Boltzmann Generators (klein2024transferable, TBG) architecture for the generative backbone model, namely TBG-CV. For time-lagged conditioning, MLCVs were concatenated to the node feature of GNNs. DeepTICA and TAE used heavy atom distances as stated, while VDE and TBG-CV used heavy atom coordinates with Kabsch alignment (kabsch1976solution) for invariance. The autocorrelation loss from VDE was applied to stabilize the training. The number of parameters of encoders were all identical, with the same training configuration under the mlcolvar library (bonati2023unified).
OPES simulation results. Meta-stable basins were defined as $\phi>0$ and $\phi<0$, using the known optimal CVs. The first 3 ns of the OPES simulations were discarded, with $\Delta F$ updated every 1 ns. In Table 11, most simulations converge close to the reference free energy difference value obtained by using dihedral angles as CVs, while TAE exhibits a relatively high variance.
Steered molecular dynamics results. We define the target hit region by dihedral angles, and steer 256 paths by only MLCVs, with $k$ searched for a hundred units. In Table 11, TBG-CV outperforms other MLCVs with high THP and low $E_{TS}$, and shows pathways crossing near the saddle points in the energy barrier in Figure 16. while baselines mostly ignore the energy landscape.
Table 11: Quantitative results of 20 ns OPES simulations and steered MD on Alanine Dipeptide. Free energy difference values within the range of 1.25kJ/mol1.25\ $\mathrm${kJ/mol} from the reference are considered to success. RMSD and THP are averaged over all trajectories, while max energy ($E_{TS}$) is averaged over trajectories only hitting the target state. Best results are highlighted in bold and second in underline in Steered MD.
Method OPES SMD
SIGMA $\Delta F$ $k$ RMSD ($\downarrow$) THP ($\uparrow$) $E_{TS}$ ($\downarrow$)
Å % $\mathrm{kJmol}^{-1}$
Ref ($\phi$, $\psi$) 0.05 10.06 $\pm$0.22 200 1.0640 100.00 00-3.89 $\pm$5.8000
DeepTICA 0.10 09.99 $\pm$0.21 400 0.9729 008.59 814.52 $\pm$115.74
TAE 0.05 09.22 $\pm$1.74 1200 1.0086 058.59 755.41 $\pm$92.300
VDE 0.05 10.11 $\pm$0.28 700 0.8582 005.08 901.69 $\pm$115.59
TBG-CV 0.05 09.83 $\pm$1.15 300 0.9593 060.93 033.58 $\pm$15.190
(a) Ref (phi, psi)
(b) DeepTICA
(c) TAE
(d) VDE
(e) TBG-CV
Figure 16: Visualization of hitting pathways from steered MD. Initial state $C5$ and target state $C7_{ax}$ are each denoted as white circles and stars, and the red circle indicates the target hit region.
C.7 High-dimensional CVs
Here, we show qualitative results of extending CVs to higher dimensions, bigger than one. We train an encoder with four-dimensional CVs on BBA, with all other configurations identical.
Figure 17: Qualitative results of multi dimension CVs on BBA. From top to bottom, we plot the secondary structure violin plot for two residues, sensitivity analysis, visualization of sensitivity in the folded structure, and TICA plots colored by the CV value.
In Figure 17, we visualize all the analyses used in the main paper for each dimension of the CVs. While all CVs are most sensitive to the 11th and 17th residue, other sensitivities differ, indicating each dimension is looking at different input descriptors and shows a different detailed slow degree of freedom. Furthermore, the secondary structures and TICA plots between the first and other dimensions differ. Overall, the first dimension shows distinct behavior compared to other dimensions.
C.8 Longer simulation results
@@
As seen in Tables 12 and 18, simulations with approximately ten times longer results do not fully converge and still exhibit uncertainty. While the folded state is sampled many times, the unfolded states are not sampled properly, resulting in a gap in the negative range of CVs in the PMF plot. Overall, longer OPES evaluations do not provide more meaningful insights.
Appendix D Ablation experiments results
D.1 Time-lagged conditioning and fixed BioEmu weights
Components OPES simulations Steered MD
Å % $\mathrm{kJ/mol}$
D.2 Encoder sizes and placement
Param. Repr. RMSD ($\downarrow$) THP ($\uparrow$) $E_{TS}$ ($\downarrow$)
Å % $\mathrm{kJ/mol}$
48K single 1.20 $\pm$0.33 100.0 -82055.15 $\pm$098.48
1.27M single 1.50 $\pm$0.30 100.0 -82071.13 $\pm$100.08
196K single 1.41 $\pm$0.32 093.8 -82049.40 $\pm$098.52
48K pair 1.79 $\pm$0.45 056.3 -82088.33 $\pm$76.61
48K single, pair 1.66 $\pm$0.56 068.8 -82086.61 $\pm$74.80
Encoder size. We test 48K, 196K, 1.27M parameters with layers {2, 4, 8} and hidden dimension {100, 200, 400} for Chignolin, where the main paper encoder corresponds to the 48K parameter setting. All MLCVs were mixed with $C_{\alpha}$ RMSD CVs at a 5:5 ratio. In Figures 19 and 14, performance remains largely unchanged across the encoder size. Since the encoder is inferred millions of times during the OPES simulation, keeping a relatively smaller size would be practical.
D.3 Steered MD without mixing CVs
Å % $\mathrm{kJ/mol}$
Chignolin 1000 DeepTICA 07.24 $\pm$1.49 0.0 N/A
TAE 05.72 $\pm$1.94 6.2 -81951.53 $\pm$0.00
VDE N/A N/A N/A
BioEmu-CV 07.82 $\pm$1.47 0.0 N/A
2000 DeepTICA 07.01 $\pm$1.71 0.0 N/A
TAE 06.39 $\pm$1.74 0.0 N/A
VDE N/A N/A N/A
BioEmu-CV 06.09 $\pm$2.21 6.2 -82023.05 $\pm$0.00
Trp-cage 2000 DeepTICA 2.23 $\pm$1.23 0.0 N/A
TAE 8.62 $\pm$3.25 0.0 N/A
VDE N/A N/A N/A
BioEmu-CV 12.17 $\pm$1.58 0.0 N/A
5000 DeepTICA 10.78 $\pm$1.15 0.0 N/A
TAE 7.22 $\pm$1.71 0.0 N/A
VDE N/A N/A N/A
BioEmu-CV 7.14 $\pm$1.52 0.0 N/A
BBA 50000 DeepTICA 07.39 $\pm$2.71 0.0 N/A
TAE 09.87 $\pm$1.59 0.0 N/A
VDE N/A N/A N/A
BioEmu-CV 06.03 $\pm$2.56 0.0 N/A
100000 DeepTICA N/A N/A N/A
TAE 10.20 $\pm$1.59 0.0 N/A
VDE N/A N/A N/A
BioEmu-CV 06.64 $\pm$1.95 0.0 N/A

### v3

@@ 5 Conclusion
We present BioEmu-CV, a lightweight framework that learns collective variables (CVs) from time-lagged conditioning signals on a frozen BioEmu foundation model. Given the current structure, we derive a low-dimensional CV and inject it into BioEmu’s single representation, where only this conditioning path is trained with a denoising score-matching objective, leaving the backbone untouched. We systematically evaluate MLCVs on three fast-folding proteins for two downstream tasks, estimating free-energy differences using OPES simulations and transition-path sampling using CV-steered MD. Our experiments show that BioEmu-CV successfully identifies important macroscopic movements related to the slow degree of freedom, and our qualitative evaluation of state discrimination corroborates this.
@@
All protein trajectory data used in this study is publicly available from the D. E. Shaw Research (DESRES) database. The detailed description of the dataset can be found in lindorff2011fast. We provided brief explanation of the dataset in Section A.2. Implementation details for our method and all baselines, including model and training hyperparameters, are specified in Section C.2. Furthermore, the experimental setups for our downstream evaluation tasks are provided in the appendix; Section C.4 contains the configurations for the on-the-fly probability enhanced sampling (OPES) simulations, and Section C.6 outlines the setup for the steered molecular dynamics (SMD) simulations. For reproduction, we provide our code for training & simulation, and simulation environment set up.
Acknowledgments
This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Ministry of Science and ICT(MSIT) (No. RS-2022-NR072184), Korea Health Industry Development Institute(KHIDI) grant funded by the Ministry of Health & Welfare(MOHW) (No. N0425208, Developing a Highly Multimodal (Drug, Protein, Gene, Cell Imaging, Literature) Foundation Model for ADMET Property Prediction), and the Institute for Information & communications Technology Planning & Evaluation(IITP)grant funded by the Korea government(MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Program(KAIST)).
@@ A.3 Time-lagged independent component analysis
Appendix B Ablation experiments results
B.1 Time-lagged conditioning and fixed BioEmu weights
Components OPES simulations Steered MD
Å % $\mathrm{kJ/mol}$
B.2 Encoder sizes and placement
Param. Repr. RMSD ($\downarrow$) THP ($\uparrow$) $E_{TS}$ ($\downarrow$)
Å % $\mathrm{kJ/mol}$
48K single 1.20 $\pm$0.33 100.0 -82055.15 $\pm$098.48
1.27M single 1.50 $\pm$0.30 100.0 -82071.13 $\pm$100.08
196K single 1.41 $\pm$0.32 093.8 -82049.40 $\pm$098.52
48K pair 1.79 $\pm$0.45 056.3 -82088.33 $\pm$76.61
48K single, pair 1.66 $\pm$0.56 068.8 -82086.61 $\pm$74.80
Encoder size. We test 48K, 196K, 1.27M parameters with layers {2, 4, 8} and hidden dimension {100, 200, 400} for Chignolin, where the main paper encoder corresponds to the 48K parameter setting. All MLCVs were mixed with $C_{\alpha}$ RMSD CVs at a 5:5 ratio. In Figures 8 and 9, performance remains largely unchanged across the encoder size. Since the encoder is inferred millions of times during the OPES simulation, keeping a relatively smaller size would be practical.
B.3 Steered MD without mixing CVs
Å % $\mathrm{kJ/mol}$
Chignolin 1000 DeepTICA 07.24 $\pm$1.49 0.0 N/A
TAE 05.72 $\pm$1.94 6.2 -81951.53 $\pm$0.00
VDE N/A N/A N/A
BioEmu-CV 07.82 $\pm$1.47 0.0 N/A
2000 DeepTICA 07.01 $\pm$1.71 0.0 N/A
TAE 06.39 $\pm$1.74 0.0 N/A
VDE N/A N/A N/A
BioEmu-CV 06.09 $\pm$2.21 6.2 -82023.05 $\pm$0.00
Trp-cage 2000 DeepTICA 2.23 $\pm$1.23 0.0 N/A
TAE 8.62 $\pm$3.25 0.0 N/A
VDE N/A N/A N/A
BioEmu-CV 12.17 $\pm$1.58 0.0 N/A
5000 DeepTICA 10.78 $\pm$1.15 0.0 N/A
TAE 7.22 $\pm$1.71 0.0 N/A
VDE N/A N/A N/A
BioEmu-CV 7.14 $\pm$1.52 0.0 N/A
BBA 50000 DeepTICA 07.39 $\pm$2.71 0.0 N/A
TAE 09.87 $\pm$1.59 0.0 N/A
VDE N/A N/A N/A
BioEmu-CV 06.03 $\pm$2.56 0.0 N/A
100000 DeepTICA N/A N/A N/A
TAE 10.20 $\pm$1.59 0.0 N/A
VDE N/A N/A N/A
BioEmu-CV 06.64 $\pm$1.95 0.0 N/A
Appendix C Simulation and evaluation details
C.1 Qualitative verification
@@ C.2 Baseline details
C.3 Simple baselines
Experiment setup. Additionally, we show that simple baselines, e.g., PCA and TICA, on $C_{\alpha}$-wise distance fail for Trp-cage. We use the first principal component and the first time-lagged principal component as CVs. All other details are identical; the training data is for the baseline and BioEmu-CV, normalization to the range [-1, 1] over the whole DESRES trajectory data.
@@ C.3 Simple baselines
C.4 OPES simulation details
OPES configurations. We use configurations as in Table 12 for PLUMED, differing only in temperatures. A PACE of 500 steps is applied, i.e., 1 picoseconds. We use the same SIGMA value 0.05 and BARRIER value 30 for all systems. Temperatures are identical to the reference simulation. The first 100 ns are discarded as equilibration time with 50 ns window unit time steps for Figures 2 and 14.
Protein PACE SIGMA BARRIER (kJ/mol) Temperature (K)
Chignolin 500 0.05 30 340
Trp-cage (TC10b) 500 0.05 30 290
BBA 500 0.05 30 325
C.5 Longer simulation results
@@
As seen in Tables 13 and 13, simulations with approximately ten times longer results do not fully converge and still exhibit uncertainty. While the folded state is sampled many times, the unfolded states are not sampled properly, resulting in a gap in the negative range of CVs in the PMF plot. Overall, longer OPES evaluations do not provide more meaningful insights.
C.6 Steered MD simulation details
Appendix D Additional results
D.1 OPES simulation results
D.2 State discrimination
D.3 Steered MD visualization
D.4 VAMP score
Method Chignolin Trp-cage BBA
VAMP-1 VAMP-2 VAMP-E VAMP-1 VAMP-2 VAMP-E VAMP-1 VAMP-2 VAMP-E
TAE 1.9686 1.9381 1.9381 1.8787 1.7722 1.7722 1.8902 1.7925 1.7925
VDE 1.9850 1.9702 1.9702 1.9947 1.9894 1.9894 1.8469 1.7172 1.7172